OpenStudy (anonymous):

Write the slope-intercept form of the equation y-1=3/4(x-3). HELP YOU GUYS. FREAKING OUT!!

OpenStudy (anonymous):

distribute the 3/4

OpenStudy (anonymous):

3/4x-9/4

OpenStudy (anonymous):

Add the one by the Y to the other side

OpenStudy (anonymous):

y=3/4x - 3 something

OpenStudy (anonymous):

Ok! *trying to process this in my head* Thanks!

OpenStudy (anonymous):

Slope int form is y=M(slope)x+B(Y intercept)

OpenStudy (texaschic101):

y - 1 = 3/4(x - 3) -- distribute through the parenthesis y - 1 = 3/4x - 9/4 -- add 1 to both sides) y = 3/4x - 9/4 + 1 y = 3/4x - 9/4 + 4/4 y = 3/4x - 5/4

OpenStudy (anonymous):

I am getting A Lot better at this! You guys are so helpful!!

OpenStudy (texaschic101):

good to hear...thank you :)